A firm has the following revenue and cost functions. TR = 60 Q – Q^2

TC = 1 /2 Q^2 +30Q + 30 Determine the quantity level at which the firm maximizes its total profit. (Hint: use marginal revenue = marginal cost rule)


1
Expert's answer
2021-02-09T07:02:33-0500

By the definition of the marginal revenue:


"MR=\dfrac{d}{dQ}(TR)=\dfrac{d}{dQ}(60Q-Q^2)=60-2Q."

By the definition the marginal cost:


"MC=\dfrac{d}{dQ}(TC)=\dfrac{d}{dQ}(0.5Q^2+30Q+30)=Q+30."

The firm maximizes its total profit when MR=MC:


"60-2Q=Q+30,""3Q=30,""Q=10."